S=n/2(a+b)
express b in terms of a,S,and n

Answer by Boreal(15235)

I assume this is S=(n/2)(a+b)
2S/n=a+b
b=(2S/n)-a
-
let a=3 b=4 and n=20
S=70
4=(140/20)-a which is 7-3